When Asinusoidal current of maximum value of 20mA flows through aresistor ,the p.d across its ends is 400V maximum .find the energy dissipated by resistor in a time of 5minutes



1
Expert's answer
2022-03-27T10:55:41-0400

The energy dissipated by resistor

"W=\frac{I_{\max}V_{\max}}{2}t"

"W=\frac{0.02\:\rm A\times 400\: V}{2}\times (5\times 60\: \rm s)=1200\: J"